OpenStudy (anonymous):

A ball is rolling down a hill at +2.0 m/s. The ball is uniformly accelerating at +4.0 m/s2. What will the ball's velocity be 3.0 s later? +13 m/s +10 m/s +14 m/s +5.3 m/s

OpenStudy (naveenbabbar):

Use formula v=u+at, gives you v=2+4*3=14m/s